OpenStudy (anonymous):

Need help please.... (a+3)^3 (^=exponent sign)

OpenStudy (raden):

u can multiply one by one of (a+3) * (a+3) * (a+3) = ....

OpenStudy (anonymous):

a3?

OpenStudy (raden):

or use the formula : (m+n)^3 = m^3 + 3*m^2*n + 3*m*n^2 + n^3

OpenStudy (anonymous):

can you check this for me? I ended up with a^3+6a^2+12a_+12

OpenStudy (raden):

it should be like this : (a+3)^3 = a^3 + 3*a^2*3 + 3*a*3^2 + 3^3 = a^3 + 9a^2 + 27a + 27

OpenStudy (anonymous):

I am probably the worst math student ever... so any help is greatly appreciated :)

OpenStudy (anonymous):

Oh, I see where I went wrong! Thank you for your help... this is only week 1, I hope I make it through to week 9. It's gonna be the death of me.

OpenStudy (raden):

you're welcome ;)